The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
A) 5 cm B) 12 cm C) 4 cm D) 11 cm

Answer by gsmani_iyer(201)


Let the length of the base = a
So the length of the one of equal sides = 4a-8 .... (Given)
So the perimeter of the triangle = a + 2(4a-8) ..(Total length of all sides)
= a+2(4a-8) = 29 cms.
= a + 8a -16 = 29
= 9a = 45 ........(by adding 16 both the sides)
So a = 45/9 = 5
So the base = 5 cms.
one of the equal sides = (4*5)-8 = 12 cms Answer.
Is this clear to you?


gsm